# Value-State Gated Attention for Mitigating Extreme-Token Phenomena in Transformers

### v2

@@ 1 Introduction
These interconnected pathologies have profound negative impacts on model performance, quantization fidelity, and interpretability. The pathological growth of residual-state norms induces numerical instability, causing divergence or stagnant learning during training (Zhai et al., 2023). Furthermore, the extreme dynamic range of activations created by these phenomena poses significant challenges for model quantization, often resulting in a substantial accuracy degradation (Bondarenko et al., 2023; Su & Yuan, 2025). Crucially, attention sinks also compromise model interpretability (Bondarenko et al., 2023; Darcet et al., 2023; Wang et al., 2024; Lappe & Giese, 2025), as attention weights no longer reliably indicate semantic importance. The systemic nature of these issues across various Transformer models underscores the need for a fundamental architectural solution.
@@ 2 Related work
Mitigating extreme-token phenomena in pretraining. Several strategies have been proposed to mitigate extreme-token phenomena during pretraining. One line of work introduces special tokens that are designed to absorb unneeded attention, such as register tokens (Darcet et al., 2023; Barbero et al., 2025; Lappe & Giese, 2025). Another group of researchers modifies the attention mechanism itself to prevent or control attention concentration. These proposals include replacing softmax with alternative functions (Ramapuram et al., 2024; Gu et al., 2024), adding a clipping function to attention scores (Bondarenko et al., 2023), or creating a learnable sink within the mechanism to channel attention (Xiao et al., 2023b; Miller, 2023; Agarwal et al., 2025). A third approach uses gating to control a token’s contribution. While gating has been widely used in LSTMs (Hochreiter & Schmidhuber, 1997) and GRUs (Dey & Salem, 2017), as well as in Transformers(e.g.Yang et al. (2024); Gao et al. (2025)), its application to resolve optimization pathologies remains less explored. Within this domain, prior approaches (Bondarenko et al., 2023; Qiu et al., 2025) have focused on learning the predictive gating function from the input embeddings. While prior methods have focused on providing static sinks for extraneous attention or predictive gates based on input features, our work is the first to propose a fully reactive, self-regulatory gate that directly addresses the emergent optimization dynamics of the value-state.
@@ 4 Experiments
Extensive experiments and comparisons are conducted to demonstrate the advantage of VGA in various settings. We begin in Section 4.1 by validating VGA on a specifically designed synthetic task as proposed by Guo et al. (2024). Then we evaluate VGA on several language models in Section 4.2, including BERT and OPT—following the practice of  Bondarenko et al. (2023), as well as GPT-2—a seminal autoregressive model whose design became the foundation for numerous subsequent large-scale commercial models, and the evidence from which is highly instructive for training large-scale commercial LLMs (Allen-Zhu, 2024). We then present quantization results for BERT and OPT in Section 4.3. We conclude in Section 4.4 by exploring VGA’s utility in fine-tuning, specifically testing whether retrofitting the module can retrospectively cure extreme-token phenomena in existing pre-trained models.
@@ 4.1 Empirical Validation on the Bigram-Backcopy Task
Table 1: Comparison of our proposed VGA against several baselines on BERT (Devlin et al., 2019), OPT-125m (Zhang et al., 2022), and GPT-2 (Radford et al., 2019) models. We evaluate task performance using Perplexity and model stability by measuring the Max I+O Norm and Avg. kurtosis to quantify activation outliers. The best result for each metric is marked in bold. VGA consistently delivers top-tier perplexity while suppressing extreme values, demonstrating its dual benefit of boosting performance and improving model stability.
@@
  GPT-2   Vanilla   $17.03^{\pm 0.00}$   $256.84^{\pm 18.77}$   $13412.71^{\pm 3,528.30}$
  Register Tokens   $\mathbf{16.26^{\pm 0.01}}$   $73.81^{\pm 5.86}$   $4871.96^{\pm 803.12}$
  Learnable Sink   $16.51^{\pm 0.00}$   $14.50^{\pm 2.27}$   $71.08^{\pm 6.32}$
  IGA   $16.62^{\pm 0.01}$   $17.68^{\pm 4.41}$   $33.51^{\pm 5.29}$
  VGA   $16.52^{\pm 0.01}$   $\mathbf{12.37^{\pm 1.28}}$   $\mathbf{31.27^{\pm 6.71}}$
@@ 4.2 Results on Representative Language Models
Models and Datasets. Our evaluation includes three representative language models: BERT (Devlin et al., 2019), OPT-125m (Zhang et al., 2022), and GPT-2 (Radford et al., 2019). Following previous work (Bondarenko et al., 2023), we evaluate BERT and OPT-125m on a combined dataset of BookCorpus (Zhu et al., 2015) and English Wikipedia (Guo et al., 2020). Our GPT-2 implementation follows Karpathy (2022) and incorporates Rotary Position Embeddings (RoPE) (Su et al., 2024), a widely adopted technique in modern Transformer architectures. We train the 124M parameter variant [^1] The training of GPT-2 (124M) requires 8 $\times$ A100 40GB for about 5 days. Training larger variants of GPT-2 required resources beyond the scope of this study. of GPT-2 on the OpenWebText (Gao et al., 2020) dataset.
@@
Baselines. Our chosen baselines are representative of the primary competing paradigms for mitigating extreme-token phenomena, as categorized in our review of related work, allowing for a direct comparison of these distinct approaches. The vanilla softmax attention mechanism serves as our primary benchmark (Vaswani et al., 2017). We also evaluate against Register Tokens (Darcet et al., 2023; Lappe & Giese, 2025), a method that prepends a set of learnable non-semantic tokens as alternative targets for attention heads. Furthermore, we compare with Learnable Sink (Xiao et al., 2023b; Agarwal et al., 2025), which introduces a dedicated token to absorb superfluous attention scores and stabilize attention patterns—a method that serves as a representative for the broader class of strategies on providing an escape from the softmax constraint (Ramapuram et al., 2024; Gu et al., 2024; Bondarenko et al., 2023). Finally, we consider IGA (Bondarenko et al., 2023), representing prior methods that learn the gating function from the input embeddings.
@@
4.4 Using VGA for Fine-tuning Pre-trained Models
Table 3: VGA Retrofitting Analysis on Pre-trained GPT-2. Comparison of a vanilla pre-trained model against a VGA equipped pre-trained model, both with 600K iterations, and two fine-tuning strategies with 6K iterations. Key Finding: Fine-tuning with VGA, consuming only $\sim$1% of the pretraining computation, significantly rectifies the pathology to a state comparable to scratch pre-training.
  GPT-2   Vanilla   $17.03^{\pm 0.00}$   $256.84^{\pm 18.77}$   $13412.71^{\pm 3,528.30}$
  VGA   $\mathbf{16.52^{\pm 0.01}}$   $\mathbf{12.37^{\pm 1.28}}$   $\textbf{31.27}^{\pm 6.71}$
  Fine-tuning w/o VGA   $17.01^{\pm 0.01}$   $398.67^{\pm 43.88}$   $9863.87^{\pm 2575.90}$
  Fine-tuning w VGA   $16.94^{\pm 0.01}$   $24.05^{\pm 4.73}$   $32.47^{\pm 2.95}$}
We investigated whether VGA could not only prevent extreme-token phenomena but also rectify them in an existing, pathologically-developed model—a concept known as retrofitting.
We performed a short fine-tuning run (6,000 steps) of a GPT-2 model on the data that was used in pre-training, comparing four variants: a vanilla pre-trained model, a VGA equipped pre-trained model, the model fine-tuned without modification (Baseline) and the model fine-tuned after retrofitting the VGA module (Retrofit).
The results in Table 3 decisively confirm VGA’s retrospective correction capability. Standard fine-tuning (w/o VGA) failed to correct the pathology; in fact, the Max I+O Norm dramatically worsened from 256.84 to 398.67, confirming the self-amplifying nature of the optimization pathology when the VGA gate is absent. In stark contrast, retrofitting the VGA module immediately stabilized the model: Max I+O Norm plummeted from 256.84 to $\mathbf{24.05}$ and Kurtosis dropped to $\mathbf{32.47}$. This strong regularization yielded the best Perplexity in the fine-tuning group ($\mathbf{16.94}$). Crucially, these stabilization metrics effectively approach to those achieved by pre-training VGA from scratch, proving that VGA is highly effective even when introduced late in the learning process, significantly expanding its practical utility.

### v3

@@ 1 Introduction
These interconnected pathologies have profound negative impacts on model performance, quantization fidelity, and interpretability. The pathological growth of residual-state norms induces numerical instability, causing divergence or stagnant learning during training (Zhai et al., 2023). Furthermore, the extreme dynamic range of activations created by these phenomena poses significant challenges for model quantization, often resulting in a substantial accuracy degradation (Bondarenko et al., 2023; Su and Yuan, 2025). Crucially, attention sinks also compromise model interpretability (Bondarenko et al., 2023; Darcet et al., 2023; Wang et al., 2024; Lappe and Giese, 2025), as attention weights no longer reliably indicate semantic importance. The systemic nature of these issues across various Transformer models underscores the need for a fundamental architectural solution.
@@ 2 Related work
Mitigating extreme-token phenomena in pretraining. Several strategies have been proposed to mitigate extreme-token phenomena during pretraining. One line of work introduces special tokens that are designed to absorb unneeded attention, such as register tokens (Darcet et al., 2023; Barbero et al., 2025; Lappe and Giese, 2025). Another group of researchers modifies the attention mechanism itself to prevent or control attention concentration. These proposals include replacing softmax with alternative functions (Ramapuram et al., 2024; Gu et al., 2024), adding a clipping function to attention scores (Bondarenko et al., 2023), or creating a learnable sink within the mechanism to channel attention (Xiao et al., 2023b; Miller, 2023; Agarwal et al., 2025). A third approach uses gating to control a token’s contribution. While gating has been widely used in LSTMs (Hochreiter and Schmidhuber, 1997) and GRUs (Dey and Salem, 2017), as well as in Transformers(e.g.Yang et al. (2024); Gao et al. (2025)), its application to resolve optimization pathologies remains less explored. Within this domain, prior approaches (Bondarenko et al., 2023; Qiu et al., 2025) have focused on learning the predictive gating function from the input embeddings. While prior methods have focused on providing static sinks for extraneous attention or predictive gates based on input features, our work is the first to propose a fully reactive, self-regulatory gate that directly addresses the emergent optimization dynamics of the value-state.
@@ 4 Experiments
Extensive experiments and comparisons are conducted to demonstrate the advantage of VGA in various settings. We begin in Section 4.1 by validating VGA on a specifically designed synthetic task as proposed by Guo et al. (2024). Then we evaluate VGA on several language models in Section 4.2, including BERT and OPT—following the practice of  Bondarenko et al. (2023). We then present quantization results for BERT and OPT in Section 4.3.
@@ 4.1 Empirical Validation on the Bigram-Backcopy Task
Table 1: Comparison of our proposed VGA against several baselines on BERT (Devlin et al., 2019), OPT-125m (Zhang et al., 2022) models. We evaluate task performance using Perplexity and model stability by measuring the Max I+O Norm and Avg. kurtosis to quantify activation outliers. The best result for each metric is marked in bold. VGA consistently delivers top-tier perplexity while suppressing extreme values, demonstrating its dual benefit of boosting performance and improving model stability.
@@ 4.2 Results on Representative Language Models
Models and Datasets. Our evaluation includes three representative language models: BERT (Devlin et al., 2019), OPT-125m (Zhang et al., 2022). Following previous work (Bondarenko et al., 2023), we evaluate BERT and OPT-125m on a combined dataset of BookCorpus (Zhu et al., 2015) and English Wikipedia (Guo et al., 2020).
@@
Baselines. Our chosen baselines are representative of the primary competing paradigms for mitigating extreme-token phenomena, as categorized in our review of related work, allowing for a direct comparison of these distinct approaches. The vanilla softmax attention mechanism serves as our primary benchmark (Vaswani et al., 2017). We also evaluate against Register Tokens (Darcet et al., 2023; Lappe and Giese, 2025), a method that prepends a set of learnable non-semantic tokens as alternative targets for attention heads. Furthermore, we compare with Learnable Sink (Xiao et al., 2023b; Agarwal et al., 2025), which introduces a dedicated token to absorb superfluous attention scores and stabilize attention patterns—a method that serves as a representative for the broader class of strategies on providing an escape from the softmax constraint (Ramapuram et al., 2024; Gu et al., 2024; Bondarenko et al., 2023). Finally, we consider IGA (Bondarenko et al., 2023), representing prior methods that learn the gating function from the input embeddings.